# Constant-roll inflation with tachyon field in the holographic braneworld

### v1

@@ 4.2 Scalar spectral index and tensor-to-scalar ratio
Figure 7: $r$ versus $n_{\rm s}$ diagrams, in the models with $\eta={\rm const}$ (left panel) and $\bar{\eta}={\rm const}$ (right panel), with observational constraints from Planck Collaboration 2018 [25]. The dots are obtained for randomly chosen $\eta$, $h_{\rm i}$ and $h_{,\theta\rm i}$ in the intervals $0.08\leq\eta\leq 0.085$, $1.39\leq h_{\rm i}\leq 1.41$ and $-0.01\leq h_{,\theta\rm i}\leq-0.9$ (left panel), and in the intervals $-0.001\leq\eta\leq-0.04$, $1.38\leq h_{\rm i}\leq 1.414$ and $-0.1\leq h_{,\theta\rm i}\leq-0.9$ (right panel). The colors represent different $N$: $N=55$ (red), $N=60$ (blue), and $N=65$ (green).
@@ 5 The swampland criteria and the braneworld inflation
First, consider the second criterion, using the parameter $\eta$ as presented in (12). The second criterion is in strong tension with the single-field slow-roll inflation model because the slow-roll parameter
$\varepsilon_{1}=\frac{M_{\rm Pl}^{2}}{2}\frac{V_{,\theta}^{2}}{V^{2}}$ (41)
should be smaller than unity to have accelerated expansion phase. The definition (41) is only valid in the slow-roll approximation, in a model with a canonical scalar field in the standard cosmology [32]. For tachyon field the therm $V_{,\theta}/V$ can be calculated using the equation for $\theta(t)$ [23]
@@
$\frac{\ell V_{,\theta}}{V}=-\frac{(1-\frac{h^{2}}{2})\left(-36h^{3}(1-\frac{h^{2}}{4})-8hh_{,\theta}^{2}+8(1-\frac{h^{2}}{2})h_{,\theta\theta}\right)h_{,\theta}}{18h^{4}(1-\frac{h^{2}}{4})^{2}-8(1-\frac{h^{2}}{2})^{2}h_{,\theta}^{2}}.$ (47)
From Fig. 8, we can conclude that our holographic model provides the second swampland criterion to be satisfied during inflation for a wide range of the relevant parameters. For instance, for $\bar{\eta}=\rm{const.}$ and $N=65$, the following set of observational parameters is obtained: $n_{\rm s}=0.964$, $r=0.055$, that are in a good agreement with the current observational data. We can roughly estimate that the second swampland criterion is satisfied for $n_{\rm s}>0.96$ and $r>0.05$. A similar behavior for $V_{,\theta}/V$ is obtained in Ref. [5].

### v2

@@ 4.5 Comparison with the Planck data
Figure 7: $r$ versus $n_{\rm s}$ diagrams, in the models with $\eta={\rm const}$, with observational constraints from Planck Collaboration 2018 [25]. The dots are obtained for randomly chosen $\eta$, $h_{\rm i}$ and $h_{,\theta\rm i}$ in the intervals $0.08\leq\eta\leq 0.085$, $1.39\leq h_{\rm i}\leq 1.41$ and $-0.01\leq h_{,\theta\rm i}\leq-0.9$. The colors represent different $N$: $N=55$ (red), $N=60$ (blue), and $N=65$ (green).
Figure 8: $r$ versus $n_{\rm s}$ diagrams, in the models with $\bar{\eta}={\rm const}$, with observational constraints from Planck Collaboration 2018 [25]. The dots are obtained for randomly chosen $\bar{\eta}$, $h_{\rm i}$ and $h_{,\theta\rm i}$ in the intervals $-0.001\leq\bar{\eta}\leq-0.04$, $1.38\leq h_{\rm i}\leq 1.414$ and $-0.1\leq h_{,\theta\rm i}\leq-0.9$. The colors represent different $N$: $N=55$ (red), $N=60$ (blue), and $N=65$ (green). The analytical results obtained in Ref. [16] are depicted by the full lines.
@@
Unfortunately, the tachyon inflation suffers from a reheating problem imminent for all tachyon models with the ground state at $\theta\rightarrow\infty$ [42]. After the inflationary epoch, the tachyon will remain a dominant component unless at the end of inflation, it decayed into inhomogeneous fluctuations and other particles. This period, known as reheating [43], links the inflationary epoch with the subsequent thermalized radiation era. Unlike in the canonical models of inflation, the tachyon field rolls towards its ground state without oscillating about it, and the conventional reheating mechanism does not work. However, it was shown [44] that coupling of massless fields to the time-dependent tachyon condensate could yield a reheating efficient enough to overcome the problem of a cold dark-matter dominance.
The authors of Ref. [45] have explicitly studied the reheating that results from a coupling of the tachyon with a U(1) gauge field through a mechanism of cosmological production. Unfortunately, it turns out that the proposed scenario alone cannot solve the reheating problem of the tachyon inflation. It would be of considerable interest to investigate the effects of cosmological creation in the warm inflation models [46]. In warm inflation, radiation due to dissipative effects is produced in parallel with the inflationary expansion, and inflation ends when the universe heats up to become radiation-dominated. This scenario has been successfully applied to tachyon inflation models [47] and, in principle, should also work for tachyon inflation in the holographic braneworld inflation presented here. However, this study is beyond the scope of the present paper.
@@ 5 The swampland criteria and the braneworld inflation
Consider the second criterion. In the single-field slow-roll inflation model the first slow-roll parameter is given by
$\varepsilon_{1}=\frac{M_{\rm Pl}^{2}}{2}\frac{V_{,\theta}^{2}}{V^{2}},$ (83)
and the second swampland criterion implays $\varepsilon_{1}>c^{\prime 2}/2$ [49]. Note that for $c^{\prime}\sim 0.1$, one obtains a consistency of the model with both the observational data and the swampland conjecture. Namely, as a consequence of (82) and (83), to satisfy the second swampland criteria one needs $\varepsilon_{1}>0.005$, in perfect agreement with the condition $\varepsilon_{1}<1$ for accelerated expansion.
Next, we demonstrate that in our holographic approach the parameter $\eta=\rm{const}$ is not consistent with the second swampland criterion, whereas $\bar{\eta}=\rm{const}$ may satisfies both the second swampland criterion and Planck data. The definition (83) is only valid in the slow-roll approximation, in a model with a canonical scalar field in the standard cosmology [32]. For tachyon field the therm $V_{,\theta}/V$ can be calculated using the equation for $\theta(t)$ [23]
@@
$\frac{\ell V_{,\theta}}{V}=-\frac{(1-\frac{h^{2}}{2})\left(-36h^{3}(1-\frac{h^{2}}{4})-8\ell^{2}hh_{,\theta}^{2}+8\ell^{2}(1-\frac{h^{2}}{2})h_{,\theta\theta}\right)\ell h_{,\theta}}{18h^{4}(1-\frac{h^{2}}{4})^{2}-8\ell^{2}(1-\frac{h^{2}}{2})^{2}h_{,\theta}^{2}}.$ (89)
From Fig. 9, we can conclude that our holographic model provides the second swampland criterion to be satisfied during whole inflation for a wide range of the relevant parameters. For instance, for $\bar{\eta}=\rm{const.}$ and $N=65$, the following set of observational parameters is obtained: $n_{\rm s}=0.964$, $r=0.055$, that are in a good agreement with the current observational data. We can roughly estimate that the second swampland criterion is satisfied for $n_{\rm s}>0.96$ and $r>0.05$. A similar behavior for $V_{,\theta}/V$ is obtained in Ref.  [5].